# A Neurosymbolic Approach to Adaptive Feature Extraction in SLAM

### v2

@@ Abstract
Autonomous robots, autonomous vehicles, and humans wearing mixed-reality headsets require accurate and reliable tracking services for safety-critical applications in dynamically changing real-world environments. However, the existing tracking approaches, such as Simultaneous Localization and Mapping (SLAM), do not adapt well to environmental changes and boundary conditions despite extensive manual tuning. On the other hand, while deep learning-based approaches can better adapt to environmental changes, they typically demand substantial data for training and often lack flexibility in adapting to new domains. To solve this problem, we propose leveraging the neurosymbolic program synthesis approach to construct adaptable SLAM pipelines that integrate the domain knowledge from traditional SLAM approaches while leveraging data to learn complex relationships. While the approach can synthesize end-to-end SLAM pipelines, we focus on synthesizing the feature extraction module. We first devise a domain-specific language (DSL) that can encapsulate domain knowledge on the important attributes for feature extraction and the real-world performance of various feature extractors. Our neurosymbolic architecture then undertakes adaptive feature extraction, optimizing parameters via learning while employing symbolic reasoning to select the most suitable feature extractor. Our evaluations demonstrate that our approach, neurosymbolic Feature EXtraction (nFEX), yields higher-quality features. It also reduces the pose error observed for the state-of-the-art baseline feature extractors ORB and SIFT by up to 90% and up to 66%, respectively, thereby enhancing the system’s efficiency and adaptability to novel environments.
@@ I Introduction
In this work, we formalize these hybrid compositional approaches as neurosymbolic programs [9, 10, 11] which aim to bridge the gap between data-driven learning approaches and rule-based symbolic reasoning. In particular, formulating the SLAM pipeline as a composition of modules–each potentially represented as a neurosymbolic program–enables neurosymbolic program synthesis, where we aim to synthesize tracking programs given a library of neural and symbolic components that fit a dataset and generalize to unseen inputs [9]. While neurosymbolic program synthesis can be employed at the module level or the entire tracking pipeline, we focus on the feature extraction module of the SLAM pipeline as a proof of concept. The feature extraction module detects and tracks feature maps over time. There has been significant prior work on developing feature extraction approaches, such as ORB [12] and SIFT [13]. These techniques are well-suited for certain environmental conditions for specific applications [14, 15]. Therefore, the feature extraction performance can be improved by dynamically selecting the most appropriate feature extractor and adapting its configuration parameters for the given environmental scenario. However, the search space across feature extractors, configuration parameters, agent types, and scenarios is large, and extensive performance data across all possible combinations may be limited.
@@
In this work, we consider finding the right feature extractor and its parameters a neurosymbolic program synthesis task. The neural component of our neurosymbolic feature extractor, nFEX, is a standard neural network that finds the optimal parameters for the various feature extractors we consider using a dataset of optimal configurations under different environmental conditions. We empirically generate this dataset by running exhaustive combinations of feature extractors and their parameters through SLAM pipelines. nFEX’s symbolic component captures the domain knowledge on the attributes of essential features and the impact of environmental conditions on the end-to-end performance of a SLAM pipeline as well as the feature quality metrics, such as texture, dissimilarity, and spatial density. Moreover, the symbolic representation can incorporate prior knowledge on the end-to-end performance of feature extractors, e.g., ORB gives the lowest pose error under bright conditions with good textures. As with any neurosymbolic program synthesis task, we specify the symbolic component using the syntax of a domain-specific language (DSL) [16], and we leverage knowledge graphs to represent any symbolic expert knowledge. Given these components, we devise a learning algorithm that uses symbolic reasoning to adapt the outcomes of the neural learning component. The task for the algorithm is to discover the program’s discrete architecture (i.e., a feature extractor) and its real-valued parameters (i.e., feature extractor’s configurations). The task specification that directs this search includes a domain-specific quantitative fitness function derived from labeled data or expert knowledge. The algorithm aims to find a program that optimizes the loss under constraints.
@@
The task of neurosymbolic program synthesis is not trivial and requires solving multiple challenges. First, there are no existing SLAM pipelines or feature extraction module DSLs. Therefore, we need to devise a DSL to represent various agent types (car, drone, human), their motion types (fast, slow), scene types (indoor, outdoor), light conditions (bright, dark), the set of feature extractors (ORB, SIFT), and their parameters (e.g., no. of features, density), details in §IV. Second, we must populate a knowledge graph summarizing how different feature extraction approaches perform under various conditions. However, as such data is limited in the existing literature [14, 15], we employed an experiment-driven approach to populate the knowledge graph (details in §III-A1). We present our solution to the aforementioned and other challenges in their respective sections.
@@ II-C Neurosymbolic Architectures
Recent studies [11, 29] have highlighted the effectiveness of neurosymbolic program synthesis [30] in advancing the capabilities of machines to understand [10, 31], navigate [32], and interact with their environments [33], demonstrating a practical approach to enhancing feature extraction through the integration of learning and reasoning processes. Neurosymbolic architectures can address the main issues of usual SLAM feature extractors, which can not easily adjust to various environmental changes, affecting their accuracy and reliability. This approach uses the strength of neural networks to process complex data and symbolic systems to apply specific knowledge and rules [9], making it better suited for different conditions.
@@ IV-A2 Feature Extractor Selection ()
Once we obtain for each feature extractor , given set of symbolic metrics with dynamically adjusted weights for each . is a function that computes the weight of the metric based on the environmental conditions . We assess the quality of detected features for each feature extractor with its parameters optimized as as:
@@ IV-B3 Feature quality metrics ()
Our choice of emphasizes metrics intrinsic to the feature extraction process, ensuring our neurosymbolic optimization can directly influence them without relying on data from downstream SLAM components.
@@
IV-B4 Datasets
IV-B5 Fitness Function Training
We use a two-step neural network to optimize feature extraction parameters and feature extractor selection . training employs a Multi-Layer Perceptron (MLP), with three fully connected layers and ReLU activations, against a mean squared error (MSE) loss of the pose error. Then, training employs a hybrid neural network to evaluate feature extractor efficacy by processing numerical data and images. This stage uses Convolutional Neural Network (CNN) layers for image analysis integrated with fully connected layers for numerical insights, assessing the feature quality. The network minimizes loss function based on the cross-entropy of the feature quality scores. Feature scaling is applied in both models using a standard scaler to normalize the features, making them more conducive for neural network training by ensuring they have a mean of zero and a standard deviation of one.
To establish ground truth for optimized and feature extractor selection , we crafted 144 combinations for each with the following set: NF , SF , NL , ST .
V Evaluation
This section evaluates the integration of nFEX with ORBSLAM3; we use Absolute Trajectory Error (ATE) for performance comparison. It calculates the deviation between aligned estimated and ground truth trajectories as: . Though precise trajectories do not guarantee flawless maps, they are widely used to gauge SLAM performance. We use ORB and SIFT with default (parameters from OpenCV) and dynamically optimized parameters () as baselines, and we trained our models on sequences EuroC-MH_02, KITTI-00, and Holoset-Suburb-seq-1. Lastly, we ran all our tests on NVIDIA GeForce RTX 2070 GPU, with a batch size of 8, Adam optimizer, and a learning rate of .
@@ V-C1 Impact of Parameter Optimization
In Figure 6, we examine the effect of individual parameters— —on the ATE for the ORB feature extractor. We fixed one parameter and kept the other dynamic. ORB with all dynamic or all default parameters are highlighted by black and red lines, respectively. This comparison shows the potential of dynamic parameter configurations, particularly in complex and varied environments. Further analysis shows nFEX and dynamically optimized ORB (ORB-Dyn) significantly increase feature matches over default settings (Table III). nFEX notably outperforms default ORB (ORB-Def) and SIFT with dynamic parameters (SIFT-Dyn), showcasing its ability to adaptively optimize feature extraction for superior match quality.
@@ V-C2 Impact of Feature Extractor Selection
In Table IV, we show nFEX’s feature extractor selection capability on the sequence, which dynamically chooses between ORB and SIFT extractors. The selection process yields a varied distribution: ORB-Dyn is chosen for 1366 out of 3638 frames, showcasing its preference in certain conditions, followed by SIFT-Dyn for 1124 frames. ORB-Def and SIFT-Def (SIFT with default parameter) are less frequently selected, indicating that dynamic parameter optimization generally outperforms default settings.
@@ V-D Computation
In Table V, we show the computational efficiency of nFEX averaged over 10 runs. Model loading takes 1800 milliseconds (ms), which is a one-time overhead. Parameter and feature extractor selection are executed swiftly, ensuring rapid adaptability. At 112 ms, frame processing remains within acceptable bounds for real-time application, positioning nFEX as a viable solution for enhancing SLAM performance with negligible impact on processing speed.
@@ VII Conclusion
Accurate and adaptive tracking remains a key challenge for achieving autonomous agents’ safe and robust operation in dynamic environments. Our approach leverages the strengths of both symbolic reasoning and data-driven learning to dynamically select and configure the most appropriate feature extractor based on the encountered environment. We demonstrated how nFEX’s dynamicity significantly improves performance across three representative benchmark datasets in different domains compared to traditional approaches. While this work focuses on adaptive SLAM feature extraction, the modular nature of the SLAM pipeline paves the way for extending neurosymbolic program synthesis to other modules as well.

### v3

@@ Abstract
Autonomous robots, autonomous vehicles, and humans wearing mixed-reality headsets require accurate and reliable tracking services for safety-critical applications in dynamically changing real-world environments. However, the existing tracking approaches, such as Simultaneous Localization and Mapping (SLAM), do not adapt well to environmental changes and boundary conditions despite extensive manual tuning. On the other hand, while deep learning-based approaches can better adapt to environmental changes, they typically demand substantial data for training and often lack flexibility in adapting to new domains. To solve this problem, we propose leveraging the neurosymbolic program synthesis approach to construct adaptable SLAM pipelines that integrate the domain knowledge from traditional SLAM approaches while leveraging data to learn complex relationships. While the approach can synthesize end-to-end SLAM pipelines, we focus on synthesizing the feature extraction module. We first devise a domain-specific language (DSL) that can encapsulate domain knowledge on the essential attributes for feature extraction and the real-world performance of various feature extractors. Our neurosymbolic architecture then undertakes adaptive feature extraction, optimizing parameters via learning while employing symbolic reasoning to select the most suitable feature extractor. Our evaluations demonstrate that our approach, neurosymbolic Feature EXtraction (nFEX), yields higher-quality features. It also reduces the pose error observed for the state-of-the-art baseline feature extractors ORB and SIFT by up to 90% and up to 66%, respectively, thereby enhancing the system’s efficiency and adaptability to novel environments.
@@ I Introduction
In this work, we formalize these hybrid compositional approaches as neurosymbolic programs [9, 11], which aim to bridge the gap between data-driven learning approaches and rule-based symbolic reasoning. In particular, formulating the SLAM pipeline as a composition of modules–each potentially represented as a neurosymbolic program–enables neurosymbolic program synthesis, where we aim to synthesize tracking programs given a library of neural and symbolic components that fit a dataset and generalize to unseen inputs [9]. While neurosymbolic program synthesis can be employed at the module level or the entire tracking pipeline, we focus on the feature extraction module of the SLAM pipeline as a proof of concept. The feature extraction module detects and tracks feature maps over time. Significant prior work has been on developing feature extraction approaches, such as ORB [12] and SIFT [13]. These techniques are well-suited for certain environmental conditions for specific applications [14, 15]. Therefore, the feature extraction performance can be improved by dynamically selecting the most appropriate feature extractor and adapting its configuration parameters for the given environmental scenario. However, the search space across feature extractors, configurations, agent types, and scenarios is large, with limited performance data available for all combinations.
@@
In this work, we consider finding the right feature extractor and its parameters a neurosymbolic program synthesis task. The neural component of our neurosymbolic feature extractor, nFEX, is a standard neural network that finds the optimal parameters for the various feature extractors we consider using a dataset of optimal configurations under different environmental conditions. We empirically generate this dataset by running exhaustive combinations of feature extractors and their parameters through SLAM pipelines. nFEX’s symbolic component captures the domain knowledge on the attributes of essential features and the impact of environmental conditions on the end-to-end performance of a SLAM pipeline as well as the feature quality metrics, such as texture, dissimilarity, and spatial density. Moreover, the symbolic representation can incorporate prior knowledge on the end-to-end performance of feature extractors; for example, ORB gives the lowest pose error under bright conditions with good textures. As with any neurosymbolic program synthesis task, we specify the symbolic component using the syntax of a domain-specific language (DSL) [16], and we leverage knowledge graphs to represent any symbolic expert knowledge. Given these components, we devise a learning algorithm that uses symbolic reasoning to adapt the outcomes of the neural learning component. The task for the algorithm is to discover the program’s discrete architecture (i.e., a feature extractor) and its real-valued parameters (i.e., feature extractor’s configurations). The task specification that directs this search includes a domain-specific quantitative fitness function derived from labeled data or expert knowledge. The algorithm aims to find a program that optimizes the loss under constraints.
@@
The task of neurosymbolic program synthesis is not trivial and requires solving multiple challenges. First, there are no existing SLAM pipelines or feature extraction module DSLs. Therefore, we need to devise a DSL to represent various agent types (car, drone, human), their motion types (fast, slow), scene types (indoor, outdoor), light conditions (bright, dark), the set of feature extractors (ORB, SIFT), and their parameters (e.g., no. of features, density), details in §IV. Second, we must populate a knowledge graph summarizing how different feature extraction approaches perform under various conditions. However, as such data is limited in the existing literature [14, 15], we employed an experiment-driven approach to populate the knowledge graph (details in §III-A1). We present our solution to those mentioned above and other challenges in their respective sections.
@@ II-C Neurosymbolic Architectures
Recent studies [11, 29, 10] have highlighted the effectiveness of neurosymbolic program synthesis [30] in advancing the capabilities of machines to understand [31, 10], navigate [32], and interact with their environments [33], demonstrating a practical approach to enhancing feature extraction through the integration of learning and reasoning processes. Neurosymbolic architectures can address the main issues of usual SLAM feature extractors, which can not easily adjust to various environmental changes, affecting their accuracy and reliability. This approach uses the strength of neural networks to process complex data and symbolic systems to apply specific knowledge and rules [9], making it better suited for different conditions.
@@ IV-A2 Feature Extractor Selection ()
Once we obtain for each feature extractor , given set of symbolic metrics with dynamically adjusted weights for each . is a function that computes the weight of the metric based on the environmental conditions . We assess the quality of detected features for each feature extractor with its parameters optimized as :
@@ IV-B3 Feature quality metrics ()
Our choice of focuses on metrics intrinsic to feature extraction, allowing our neurosymbolic optimization to influence them without relying on downstream SLAM data.
@@
IV-B4 Datasets
IV-B5 Fitness Function Training
We use a two-step neural network to optimize feature extraction parameters and feature extractor selection . training uses a Multi-Layer Perceptron (MLP) fully connected layers () and ReLU activations (), minimizing the loss as:
where is the batch size, the true label, and the predicted value. Next, training uses a hybrid neural network combining Convolutional Neural Network (CNN) layers for image analysis and fully connected layers () for numerical data, refining feature quality by minimizing the cross-entropy loss, which quantifies the discrepancy between predicted probabilities and true labels as:
Feature scaling is applied using a standard scaler to ensure features have a mean of zero and a standard deviation of one, making them more suitable for neural network training.
V Evaluation
We evaluate the integration of nFEX with ORBSLAM3 by comparing performance using the Absolute Trajectory Error (ATE). It calculates the deviation between aligned estimated and ground truth trajectories as: . Here represents the total number of trajectory points being compared between the and . Though precise trajectories do not guarantee flawless maps, they are widely used to gauge SLAM performance. We use ORB and SIFT with default (parameters from OpenCV) and dynamically optimized parameters () as baselines, and we trained our models on sequences EuroC-MH_02, KITTI-00, and Holoset-Suburb-seq-1. Lastly, we ran all our tests on NVIDIA GeForce RTX 2070 GPU, with a batch size of 8, Adam optimizer, and a learning rate of .
@@ V-C1 Impact of Parameter Optimization
In Figure 6, we examine the effect of individual parameters— —on the ATE for the ORB feature extractor. We fixed one parameter and kept the other dynamic. ORB with all dynamic or all default parameters are highlighted by black and red lines, respectively. This comparison shows the potential of dynamic parameter configurations, particularly in complex and varied environments. Further analysis shows nFEX and dynamically optimized ORB (ORB-Dyn) significantly increase feature matches over default settings (Table III). nFEX notably outperforms default ORB (ORB-Def) and SIFT with dynamic parameters (SIFT-Dyn), showcasing its ability to optimize feature extraction for superior match quality adaptively.
@@ V-C2 Impact of Feature Extractor Selection
Table IV illustrates nFEX’s feature extractor selection on the sequence, which dynamically switches between ORB and SIFT extractors. The selection process results in a diverse distribution: ORB-Dyn is selected for 1366 out of 3638 frames, demonstrating its preference under specific conditions, followed by SIFT-Dyn with 1124 frames. In contrast, ORB-Def and SIFT-Def (SIFT with default parameters) are chosen less frequently, indicating that dynamic parameter optimization generally outperforms default settings.
@@ V-D Computation
In Table V, we show the computational efficiency of nFEX averaged over 10 runs. Model loading takes 1800 milliseconds (ms), a one-time overhead. Parameter and feature extractor selection are executed swiftly, ensuring rapid adaptability. At 112 ms, frame processing remains within acceptable bounds for real-time application, positioning nFEX as a viable solution for enhancing SLAM performance with negligible impact on processing speed.
@@ VII Conclusion
Accurate and adaptive tracking remains a key challenge for achieving agents’ safe and robust operation in dynamic environments. Our approach leverages the strengths of both symbolic reasoning and data-driven learning to dynamically select and configure the most appropriate feature extractor based on the encountered environment. We demonstrated how nFEX’s dynamicity significantly improves performance across three representative benchmark datasets in different domains compared to traditional approaches. While this work focuses on adaptive SLAM feature extraction, the modular nature of the SLAM pipeline paves the way for extending neurosymbolic program synthesis to other modules as well.
Acknowledgments
The research reported in this paper was sponsored in part by the National Science Foundation (NSF) under awards 2435642 and 2237485. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the funding agencies.